OpenStudy (anonymous):

evaluate the expression sqrt. 8-n when n=8

OpenStudy (anonymous):

its 0

OpenStudy (anonymous):

sqrt(8-n) when n = 8 is sqrt(8-8) = sqrt(0) = 0.

OpenStudy (anonymous):

Hey pollo, can we help you understand what it means to evaluate an expression rather than just solving problems for you?

OpenStudy (anonymous):

yeah that would be great!

OpenStudy (anonymous):

Do you have a question about how this works? All of the questions you've asked so far just ask you to put the number in for the letter.

OpenStudy (anonymous):

like in this one, n is 8.... so you just put in 8 where you see the letter n

OpenStudy (anonymous):

yes how does this work?

OpenStudy (anonymous):

so I have a different one for you real quick

OpenStudy (anonymous):

i see just replacing them in the expression

OpenStudy (anonymous):

x = 100 so what is x+50?

OpenStudy (anonymous):

put in 100 where x is first and tell me what that is.

OpenStudy (anonymous):

150

OpenStudy (anonymous):

yup :D, but show each step.

OpenStudy (anonymous):

so, 100+50 = 150

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I hope that helps you out.

OpenStudy (anonymous):

it actually makes sense thank you very much!

OpenStudy (anonymous):

no problem!